A 5000 is invested for 3 years at 10 percent,compounded annually.Calculate the amount to be received after 3 years? Also calculate the amount of interest that will be earned?
Answer by lwsshak3(11628)

A 5000 is invested for 3 years at 10 percent,compounded annually.Calculate the amount to be received after 3 years? Also calculate the amount of interest that will be earned?
***
compound interest formula: A=P(1+i)^n, p=initial amount invested, i=interest rate per period, A=amount after n periods.
For given problem:
P=5000
i=10%
n=3
..
A=5000(1+.10)^3=5000(1.1)^3=6655
6655-5000=1655
amount to be received after 3 years: 6655
amount of interest that will be earned:1655